If IC = 50 mA and IE = 52 mA, calculate the current gain of the transistor


1
Expert's answer
2020-11-16T06:40:56-0500

There are two configuration mainly for transistor current amplifier.

1) Common emitter: In common emitter the current gain "(\beta) =\frac{I_C}{I_B}"


We have "I_E=I_B+I_C"

There fore "I_B=I_E-I_C=52-50=2mA"


Hence "\beta=\frac{50}{2}=25"


2)Common base: In common base current gain"(\alpha) =\frac{I_C}{I_E}=\frac{50}{52}=0.96"


Generally "\alpha" is close to one. Common emitter is more popular and widely used due to it's large gain.